(a-2)(a+4)= 2a(a-3)
Answer by checkley75(3666)

(A-2)(A+4)=2A(A-3)
A^2+2A-8=2A^2-6A
A^2-2A^2+2A+6A-8=0
-A^2+8A-8=0
A^2-8A+8=0
USING THE QUADRATIC EQUATION WE GET:
A=(8+-SQRT[-8^2-4*1*8])/2*1
A=(8+-SQRT[64-32])/2
A=(8+-SQRT32)/2
A=(8+-5.657)/2
A=(8+5.657)/2
A=13.657/2
A=6.8285 ANSWER.
A=(8-5.657)/2
A=2.343/2
A=1.1715 ANSWER.